OpenStudy (anonymous):

Multiplying Polynomials and Monomials. (-s^2t^3) (st)

OpenStudy (anonymous):

ok... do the same thing as before... but notice that negative sign? that's actually a -1...

OpenStudy (anonymous):

[-s^((2t^3)+1)]t

OpenStudy (anonymous):

Soo , I got -s^2t^3

OpenStudy (anonymous):

Would the (st) be different terms because it doesn't have an exponent ?

OpenStudy (anonymous):

s=s^1 t=t^1 so st = s^1t^1

OpenStudy (anonymous):

try again.

OpenStudy (anonymous):

So , -s^3t^4 ?

OpenStudy (anonymous):

there you go! :)

OpenStudy (anonymous):

How do you multiply something like this : (x-5) (x+5) Would you add the varibables and then do the addition ?